OpenStudy (anonymous):

@aum Malcolm bought dog-food bowls to give away as prizes at a dog-lovers event. The plastic bowls cost $6 each and the stainless steel bowls cost $10.50 each. Malcolm bought 7 plastic bowls and spent a total of $136.50 on the prizes. How many stainless steel bowls did Malcolm buy? What would the x and y for this problem?

OpenStudy (anonymous):

@jdoe0001 @austinL @karatechopper

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

lets say x is the plastic bowl and y is the steel bowl.

TheSmartOne (thesmartone):

so 6x+10.50y=136.50

TheSmartOne (thesmartone):

We know that x=7

OpenStudy (anonymous):

6x7=42 ----> 42+10.50y=136.50 What's Y?

TheSmartOne (thesmartone):

subtract 42 on both sides. So 10.50y=94.5 Divide by 10.5 on both sides y=9

TheSmartOne (thesmartone):

@CrazyCountryGirl

OpenStudy (anonymous):

OK so... 6x7=42 42+10.50y=136.50 10.50x9=94.50 42+94.50=136.50 He bought 63 bowls?

OpenStudy (anonymous):

Woops I mean 16? @TheSmartOne

OpenStudy (anonymous):

@TheSmartOne @satellite73 @e.mccormick @jim_thompson5910

jimthompson5910 (jim_thompson5910):

How many stainless steel bowls did Malcolm buy?

jimthompson5910 (jim_thompson5910):

look back at TheSmartOne's post where x and y are defined

jimthompson5910 (jim_thompson5910):

what are x and y?

OpenStudy (anonymous):

9 and 7. Do I multiply them?

jimthompson5910 (jim_thompson5910):

x = 9 means what?

OpenStudy (anonymous):

9 is one of the types of bowls?

jimthompson5910 (jim_thompson5910):

yes go back and see how TheSmartOne defined x and y

OpenStudy (anonymous):

X as the Plastic Bowls = 7 Y as the Steel Bowls = 9

jimthompson5910 (jim_thompson5910):

so because y = 9, there are 9 steel bowls

OpenStudy (anonymous):

Oh. I thought it wanted to know how many bowls there were together. :) I see. I already knew that. The answer would be 9, right?

jimthompson5910 (jim_thompson5910):

nope, they're just asking about the steel bowls

OpenStudy (anonymous):

Thanks! :)

jimthompson5910 (jim_thompson5910):

np

TheSmartOne (thesmartone):

Thanks @jim_thompson5910 I had logged out for a while. Thanks for covering my back.